(b) A balloon filled with nitrogen has a volume of 2.5dm3 at a pressure of 1.0 atm at 25°C.What will be
the volume of balloon when it is cooled to -73 °C, assuming that the pressure remains constant?
1
Expert's answer
2020-06-10T08:03:17-0400

We know that


"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"

"V_2=\frac{P_1V_1T_2}{P_2T_1}=\frac{P_2\times2.5\times200}{P_2\times298}=1.678\ dm^3"

"Since\ P_2=P_1"